From the top of a building a ball was projected horizontally the ball strikes the ground after 5 seconds at an angle of 60 degree with horizontal calculate the speed with which the ball was projected.Explain in a simple way. Thanks ๐Ÿ™๐Ÿ™

Grade:9

2 Answers

Vamshi
4 years ago
Three particles each of mass 2kg are at corners of an equilateral triangle of side โˆš3m.if one of the particle is removed,the shift in cm is
Khimraj
4 years ago
After 5 seconds , vertical component of velocity = (1/2)gt2ย = (1/2)ร—9.8ร—5ร—5 = 122.5 m/s
ย 
if u is projection velocity along horizontal direction at top of building and resultant velocity makes angle 60ยฐ with horizontal,
ยฝgt2/u = tan60
u = 70.73 m/s